# Fully Dynamic Matching and Ordered Ruzsa-Szemerédi Graphs

### v1

@@ Abstract
Our second main contribution is a better upper bound on density of both ORS and RS graphs with linear size matchings. The previous best upper bound was due to a proof of the triangle-removal lemma from more than a decade ago due to Fox [Annals of Mathematics ’11].
@@ 1.1 Our Contributions
Unfortunately there is a huge gap between existing lower and upper bounds for $\mathsf{RS}_{n}(cn)$ (and as a result also for $\mathsf{ORS}_{n}(cn)$). To the best of our knowledge, the best known upper bound on $\mathsf{RS}_{n}(cn)$ follows from the improved triangle-removal lemma due to Fox [23] from more than a decade ago which shows $\mathsf{RS}_{n}(cn)\leq n/2^{\Omega(\log^{*}n)}=n^{1-o(1)}$ (see also [24, 2]).
Our second main result in this paper is the following upper bound.
@@
For any $c>0$, it holds that $\mathsf{ORS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$, where $\log^{(x)}$ is the iterated log function.
Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the following bound on the density of RS graphs:
Corollary 3.
For any $c>0$, it holds that $\mathsf{RS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$.
Note that $\log^{(\ell)}n\gg 2^{\Omega(\log^{*}n)}$ for any fixed $c$. As such, Corollary 3 also improves the upper bound for RS graphs after more than a decade since the work of [23].
@@ 1.2 Perspective: ORS vs RS
Summarizing the above-mentioned bounds, the state-of-the-art bounds for the values of $\mathsf{RS}_{n}(cn)$ and $\mathsf{ORS}_{n}(cn)$ can be summarized as
@@ 4.1 Random Sampling Algorithm
Algorithm 1: RandomSamplingAlgorithm$(G[U],\varepsilon,\mu)$   Parameters: $\varepsilon\in(0,1]$. $M\leftarrow\emptyset$. $M_{1},...,M_{1/(2\varepsilon)}\leftarrow\emptyset$. $S\leftarrow\emptyset$ $b\leftarrow 1$. $c=0$. while $|M|<\varepsilon\mu$ do 1       if $b>\varepsilon c\cdot\log n$ then 2             $M_{c+1}\leftarrow M$. $M\leftarrow\emptyset$. $c\leftarrow c+1$. 3      for vertex $v\in U\setminus V(M)$ in random order do 4             $S\leftarrow$ Sample $b$ vertices in $U$. if $\exists u\in S$ such that $u\in U$ and $u\notin V(M)$ and $u$ is a neighbor of $v$ then 5                   $M\leftarrow(v,u)\cup M$. 6             7      $b\leftarrow 2b$ return $M$
@@
Let us consider Section 4.1, First we prove that this algorithm terminates. We need to prove that there exists $c$ such that sampling vertices with a budget $b$ in the interval of $[\varepsilon c\cdot\log n,\varepsilon(c+1)\cdot\log n]$ outputs a matching of size at least $\varepsilon\mu$. Let us assume there are no such intervals. This means that for any integer $c\in[0,1/(2\varepsilon)]$, running the algorithm for $b$ in the interval $[\varepsilon c\cdot\log n,\varepsilon(c+1)\cdot\log n]$ outputs a matching with less than $\varepsilon\mu$ edges. Note that for each edge $e$ in $M_{1}\cup\dots\cup M_{1/(2\varepsilon)}$, at most 2 edges in the certificate matching $\mu$ have common vertices with $e$. Summing over all the $1/(2\varepsilon)$ intervals we have:
$2\sum_{c=1}^{1/(2\varepsilon)}{|M_{i}|}<2\cdot\frac{1}{2\varepsilon}\cdot\varepsilon\mu=\mu.$
@@
The fact that we could not find a matching of size $\mu$ is a contradiction with the lemma statement ensuring the existence of such matching.Therefore, we have at least the size of one of the matchings $M_{1},...,M_{1/(2\varepsilon)}$ is $\varepsilon\mu$.
@@
Now, we prove the statement of the algorithm. We assume that this algorithm terminates at $\varepsilon(c-1)\cdot\log n\leq b<\varepsilon c\cdot\log n$, for a constant $\delta$, we spend $n[b+b/2+..+1]<\delta nb\leq\delta\varepsilon c\cdot n\log n<\delta\varepsilon cn^{1+\varepsilon}$ time to find the matching $M_{c}$ such that $|M_{c}|\geq\varepsilon\mu$. Note that if it takes at most $n^{2+\varepsilon}/d$ time to find the matching of size $\varepsilon n$, then we have $n^{2+\varepsilon}/d<\delta\varepsilon cn^{1+\varepsilon}$ therefore, $d>\frac{n}{\delta\varepsilon c}$.
@@
For budget $b$, we sample vertices for all the vertices, we argue that for the vertices that are not in $M$, and after sampling $b$ vertices we could not find their neighbor then the degree of vertices out of the matching is less than $n/b$. Assuming that the degree of vertex $v$ is $d_{v}$, with probability $(1-\frac{d_{v}}{n})^{b}$ we sampled a neighbor, and this probability converges to $1/e$ for $d_{v}>O(n/b)$. Therefore in this case with a constant probability we already matched $v$. This implies that if we find the first edge of this matching at a budget of $2b$, we get that the degree of any vertex out of the matching we found until this point is less than $O(n/b)$. Therefore, the total sum of the degree of the matching is less than $O(\varepsilon\mu\cdot n/b)<O(\varepsilon\mu n/((c-1)\log n))<O(\delta\varepsilon^{2}cdn/((c-1)\log n))<O(nd)$. ∎
@@ 4.2 The Overall Algorithm via Certificates and Caching Them
Algorithm 2: Algorithm for Dynamic Model in Contracted Graphs   Parameters: $\varepsilon\in(0,1]$, $t$ the threshold of the algorithm. $G^{\prime}\leftarrow$ the contracted graph. $S_{1}\leftarrow\emptyset$. $S_{2}\leftarrow\emptyset$. while  edge updates $<t$  do 1       if There are $\varepsilon^{2}n^{\prime}$new updates then 2             $S_{1}\leftarrow S_{1}\cup$ Insertions. $M\leftarrow\emptyset$. while Proposition 4 outputs a subset $U$ that has a matching of size at least $\mu=\varepsilon n^{\prime}$  do 3                   $M^{\prime}\leftarrow$ a maximal matching found by a greedy algorithm on edges of $S_{1}\cup S_{2}$ in $U$. For each edge in this set, we check if it exists in the dynamic graph. if  $|M^{\prime}|>\varepsilon\mu$  then 4                        $M\leftarrow M^{\prime}$. 5                  else 6                         $M_{U}\leftarrow$ $RandomSamplingAlgorithm(G[U\setminus V(M^{\prime})],\frac{\varepsilon}{1-2\varepsilon},(1-2\varepsilon)\mu).$ $M\leftarrow M\cup M_{U}$. 7                  $S_{2}\leftarrow S_{2}\cup M_{U}.$ Output $M$. 8             9
@@
Note that we proved in Observation 10 that if the maximal matching size is of order $n^{\prime}$, then we do not need to output a matching in $\varepsilon^{2}n^{\prime}$ updates.
@@
Algorithm 3: Main   Parameters: $\varepsilon\in(0,1]$. $k\leftarrow\log n$. $n_{1},n_{2},\dots,n_{k}\leftarrow 2,4,\dots,2^{k}$. $t\leftarrow$ threshold of algorithm Section 4.2. $\mathcal{(}C)\leftarrow$ set of maintained $O(\log n)$ contracted graphs for each $n_{c}\in\{n_{1},\dots,n_{k}\}$ using Proposition 7. $\mathcal{P}\leftarrow$ set of maintained parallel edges between two contracted sets of vertices after each update. Run a new phase Section 4.2 for each contraction after $t$ updates. For each contraction $c$ with $c_{n}$ vertices, run Proposition 5 on $c$ after $\varepsilon^{2}n_{c}$ updates to see whether $c$ preserves the maximum matching in $G$. Find the contraction $c$ preserving the largest matching of $G$ and use the output of Section 4.2 for the corresponding contraction, to output an $(1-\varepsilon)$-approximate matching.
@@ 5 Bounding Density of (Ordered) Ruzsa-Szemerédi Graphs
In this section we prove upper bounds on the density of ORS graphs. Our main result is a proof of Theorem 2 which we present in Section 5.1. We then show that a much stronger upper bound can be proved if matchings cover more than half of vertices in Section 5.2.
@@ 5.1 Linear Matchings
The following lemma, which is one of our main tools in proving Theorem 2, shows that so long as vertex degrees are larger than a threshold, we can reduce the number of vertices rather significantly without hurting the size of a relatively large number of induced matchings.
@@
[…]
@@
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 17, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
@@
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 17). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎

### v2

@@ 1.1 Our Contributions
Unfortunately there is a huge gap between existing lower and upper bounds for $\mathsf{RS}_{n}(cn)$ (and as a result also for $\mathsf{ORS}_{n}(cn)$). The best known upper bound on $\mathsf{RS}_{n}(cn)$ for linear size matchings follows from the improved triangle-removal lemma of Fox [23] which implies $\mathsf{RS}_{n}(cn)\leq n/\log^{(\ell)}$ for $\ell=O(\log(1/c))$ where $\log^{(x)}$ is the iterated log function. We note that this result is implicit in [23] and was mentioned in the paper of [24]. To our knowledge, this upper bound does not carry over to ORS graphs (we briefly discuss this at the beginning of Section 5). Our second result is a similar upper bound for ORS albeit with a worse dependence on constant $c$.
@@
For any $c>0$, it holds that $\mathsf{ORS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$.
Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the same upper bound for $\mathsf{RS}_{n}(cn)$. Note that this implication is not a new result, but the proof is very different from the abovementioned upper bound.
@@ 1.2 Perspective: ORS vs RS
Summarizing the above-mentioned bounds, we have
@@ 4.1 Random Sampling Algorithm
[…]
@@
Let us consider Section 4.1, first we prove that this algorithm will provide a matching $M_{c}$ of size a least $\varepsilon\mu/2$ by sampling vertices with budget $b\in[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$. Let us assume there are no such intervals. This means that for any integer $c\in[0,1/(2\varepsilon)]$, running the algorithm for $b$ in the interval $[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$ outputs a matching with less than $\varepsilon\mu$ edges. Note that for each edge $e$ in $M_{1}\cup\dots\cup M_{1/(2\varepsilon)}$, at most 2 edges in the maximum matching of $G[U]$ have common vertices with $e$. Summing over all the $1/(2\varepsilon)$ intervals we have:
$22\sum_{c=1}^{1/(2\varepsilon)}{|M_{i}|}<2\cdot\frac{1}{2\varepsilon}\cdot\varepsilon\mu/2=\mu/$ 2 .
@@
The fact that we could not find a matching of size $\mu/2$ is a contradiction with the lemma statement ensuring that the size of maximum matching is at least $\mu/2$. Note that by exhausting the algorithm with budget of $n$, we will go through all the edges and output a maximal matching. Note that any maximal matching in this graph should be of size at least $\mu/2$.Therefore, we have at least the size of one of the matchings in$\{M_{1},...,M_{1/(2\varepsilon)}\}$ is $\varepsilon\mu/2$.
@@
Now, we prove the statement of the algorithm. We assume that this algorithm terminates at $b$ such that $2^{\varepsilon(c-1)\cdot\log n}\leq b<2^{\varepsilon c\cdot\log n}$, for a constant $\delta$, we spend $n[b+b/2+..+1]<\delta nb\leq\delta n\cdot 2^{\varepsilon c\log n}<\delta n^{1+\varepsilon c}$ time to find the matching $M_{c}$ such that $|M_{c}|\geq\varepsilon\mu/2$. Note that if it takes at most $n^{2+\varepsilon}/d$ time to find the matching of size $\varepsilon\mu/2$, then we have $n^{2+\varepsilon}/d<\delta n^{1+\varepsilon c}$ therefore, $d>\frac{n^{1+\varepsilon-\varepsilon c}}{\delta}\geq\frac{n^{1+\varepsilon}}{\delta}$.
@@
For budget $b$, we sample vertices for all the vertices, we argue that for the vertices that are not in $M$, and after sampling $b$ vertices we could not find their neighbor then the degree of vertices out of the matching is less than $O(n/b)$. Assuming that the degree of vertex $v$ is $d_{v}=c\cdot\frac{n}{b}$, with probability $(1-\frac{d_{v}}{n})^{b}=(1-c\cdot\frac{1}{b})^{b}\approx(1/e)^{c}$ we sampled a neighbor. Therefore in this case with a constant probability we already matched $v$. This implies that if we find the first edge of this matching at a budget of $2b$, we get that the degree of any vertex out of the matching we found until this point is less than $c\cdot\frac{n}{b}$ for any constant $c$. Picking $c$ small enough, the total sum of the degree of the matching is at most $\varepsilon\mu c\cdot\frac{n}{b}<\varepsilon\mu n^{(1-\varepsilon(c-1))}=O(nd)$. ∎
@@ 4.2 The Overall Algorithm via Certificates and Caching Them
[…]
@@
Note that we proved in Observation 9 that if the maximal matching size is of order $n^{\prime}$, then we do not need to output a matching in $\varepsilon^{2}n^{\prime}$ updates.
@@
[…]
@@ 5 Bounding Density of (Ordered) Ruzsa-Szemerédi Graphs
In this section we prove upper bounds on the density of ORS graphs. Our main result of this section is a proof of Theorem 2 which we present in Section 5.1. We then show that a much stronger upper bound can be proved if matchings cover more than half of vertices in Section 5.2.
Before presenting our proof, let us first briefly discuss how the triangle removal lemma is useful for upper bounding density of RS graphs and why it does not seem to help for upper bounding ORS. The triangle removal lemma states that so long as the number of triangles in a graph are $o(n^{3})$, then we can make the graph triangle free by removing $o(n^{2})$ of its edges. Suppose we have a bipartite RS graph with induced matchings $M_{1},\ldots,M_{k}$. Add $k$ vertices $v_{1},\ldots,v_{k}$ to this graph and connect each $v_{i}$ to all the vertices matched by $M_{i}$. Now, crucially, because each $M_{i}$ is an induced matching, each $v_{i}$ is part of exactly $|M_{i}|$ triangles, one for each edge of $M_{i}$. As such, the total number of triangles can be upper bounded by $O(kn)=O(n^{2})$ which is well within the regime that one can apply triangle removal lemma. However, because the matchings in ORS are not induced matchings of the whole graph, the number of triangles cannot simply be upper bounded by $O(n^{2})$ after adding the auxiliary vertices. In fact, our upper bound completely deviates from this approach and does not rely on the triangle removal lemma.
@@ 5.1 Linear Matchings
The following lemma, which is one of our main tools in proving Theorem 2, shows that so long as vertex degrees are larger than a threshold, we can reduce the number of vertices rather significantly without hurting the size of a relatively large number of induced matchings. Our proof of Lemma 15 builds on the techniques developed for upper bounding RS graphs with matchings of size very close to $n/4$ in [24] that we extend to ORS graphs with balanced degrees.
@@
=$\mathcal{M}=\{M_{1},\ldots,M_{t/2}\}\qquad\text{and}\qquad\mathcal{M}^{\prime}=\{M_{t/2+1},\ldots,M_{t}\}.$
@@
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 16, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
@@
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 16). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎